For the production function given by q=e^0.3tK^0.75L^0.25 where the technical coefficient is denoted by e^0.3t. Both K and L grow at the rate at 1% per annum each, compute the annual growth rate of output
1
Expert's answer
2020-11-12T17:34:59-0500

q="e^{0.3t}*K^{0.75}*L^{0.25}"

technical coefficient = e^0.3t

growth in K

"K\times * \frac{101}{100}" = 1.01K

"L\times * \frac{1.01}{100}" = 1.01L

"qc=e^{0.3t}\times(1.01K^{0.75}\times1.01L^{0.25}"

qc="e^{0.3t}*1.007491K^{0.75}*1.007491L^{0.25}"

"qc=e^{0.3t}\times 1.015037K^{0.75}L^{0.25}"

annual change = "[\frac{(qc – q)}q]\times100"

annual change = "e^{0.3t}\times1.015037K^{0.75}L^{0.25} - e^{0.3t}\times K^{0.75}L^{0.25}/ e^{0.3t}K^{0.75}L^{0.25}"

annual change = "\frac{e^{0.3t}\times K^{0.75}L^{0.25}(1.015037-1)}{e^{0.3t} \times K^{0.75}L^{0.25}(1)}"

annual change = "0.015037\times100 = 1.50\%"